Calculate the rate of heat conduction through house walls that are 14.7 cm thick and that have an average thermal conductivity of 0.057 J/(s•m•ºC). Assume there are no windows or doors. The surface area of the walls is 136 m2 and their inside surface is at 19.9ºC , while their outside surface is at 4.9ºC .
Answer #1

ctt A-131 m 2 o 147
Answer #2

SOLUTION :



Rate of Heat transfer through wall

= k * (A / d) * ∆ T 


where,


k = heat transfer constant though thick surface = 0.057 J/(s*m*ºC)


A = surface area  = 136 m^2 


d = thickness = 14.7 cm = 0.147 m 


∆T = Ti - T0 = 19.9 - 4.9 = 15ºC 


So, Rate of Heat transfer = 0.057 * (136 / 0.147) * 15 = 791.02 J / s (ANSWER).

answered by: Tulsiram Garg